# FastFLUX: Pruning FLUX with Block-wise Replacement and Sandwich Training

### v1

@@ 1 Introduction
Recent advances in diffusion-based generative modelsjiang2025t2ir1reinforcingimagegeneration ; esser2024scaling ; podell2024sdxl ; sauer2023adversarialdiffusiondistillation ; chen2023pixartalphafasttrainingdiffusion ; zhang2023addingconditionalcontroltexttoimage ; zhang2024attentioncalibrationdisentangledtexttoimage ; chatterjee2024gettingrightimprovingspatial have led to unprecedented progress in high-fidelity image synthesis. Among these, transformer-based architectures such as DiT peebles2022dit and FLUX flux2024 have demonstrated strong generative capabilities by stacking deep and complex hierarchical blocks. These models often rely on extensive depth and repeated structural modules to enhance expressiveness and scalability. However, such architectural designs come at a cost: high computational complexity, large memory footprints, and latency issues that hinder real-time applications and deployment on resource-constrained platforms.
Although prior studiesli2025svdquant ; sanh2020distilbertdistilledversionbert ; meng2023distillationguideddiffusionmodels ; sun2023acceleratingdiffusionsamplingclassifierbased ; kohler2024imagineflashacceleratingemu ; dao2022flashattentionfastmemoryefficientexact ; dao2023flashattention2fasterattentionbetter ; beltagy2020longformerlongdocumenttransformer ; zaheer2021bigbirdtransformerslonger ; salimans2022progressivedistillationfastsampling have investigated model compression through pruning, knowledge distillation, and lightweight network design, these approaches typically regard the model as a black box or focus solely on parameter-level sparsity. Few methodsli2025diffusionsteprealworldsuperresolution explicitly target hierarchical redundancy, an underexplored and distinct source of inefficiency in deep generative transformers. In addition, many existing techniqueshu2023icepick ; pavlitska2024iterativefilterpruningconcatenationbased ; 10.5555/3524938.3525805 ; verdenius2020pruningiterativerankingsensitivity ; sun2024simpleeffectivepruningapproach ; 10.3844/jcssp.2025.1113.1128 ; liu2025adaptivelargelanguagemodels ; ding2025revisitinglargelanguagemodel ; frantar2023sparsegptmassivelanguagemodels require training from scratch or result in significant performance degradation, which restricts their applicability in practical deployment scenarios.
@@
In this work, we revisit the challenge of structural redundancy in deep generative models and propose FastFLUX, a general and efficient framework for architecture-level compression. It operates by selectively replacing redundant and computationally heavy residual blocks with lightweight linear alternatives, while maintaining generation quality through an efficient fine-tuning strategy. Specifically, we introduce the Block-wise Replacement with Linear Layers (BRLL) method, which targets the high-parameter residual branches in ResBlocks. In FLUX, each ResBlock consists of a heavy residual branch and a lightweight shortcut branch. We observe that the residual branches dominate computational cost and significantly slow down inference. BRLL replaces these branches with simple linear layers while preserving the original shortcut connections to retain critical information flow and ensure model stability. These replacements can be initialized via least squares estimation, allowing for training-free deployment with controllable performance degradation. To further reduce performance drop after replacement, we propose a Sandwich Training (ST) strategy. Rather than fine-tuning the entire model, ST performs localized adaptation for each replaced block by leveraging information from adjacent unpruned layers. During the training of a specific replaced block, we apply LoRAhu2021LoRAlowrankadaptationlarge modules to its neighboring layers, forming a “sandwich” structure around the linear module. This localized supervision enables the linear layer to adapt more effectively within its context, leading to better stability and quality with minimal training overhead. Using this approach, we can prune 20% of the FLUX model while maintaining high-quality image generation. The resulting model achieves significantly faster inference with only minimal performance degradation.
To sum up, the contributions of this paper are:
@@
  • We introduce Block-wise Replacement with Linear Layers (BRLL), which replaces complex residual branches with linear layers. These replacements support least squares initialization, enabling training-free deployment with controlled performance loss.
  • We propose Sandwich Training (ST) strategy to compensate for performance loss after pruning. This method performs localized fine-tuning by applying LoRA modules to the neighboring layers of each replaced block, enabling efficient and targeted adaptation with minimal training overhead.
@@ 2 Related Work
Diffusion-based Generative Models. Recent advances in diffusion models have led to impressive progress in image generation and editing. Approaches such as DDPM ho2020ddpm , Latent Diffusion rombach2022ldm , and DiT peebles2022dit demonstrate strong scalability by employing deeply stacked hierarchical modules. While these architectures enhance generative capacity, they also introduce substantial computational overhead. More recent models like FLUX flux2024 further deepen the Transformer backbone to boost fidelity, which underscores the growing need for structural efficiency and optimization in diffusion-based generation systems. This motivates a closer investigation into architectural redundancy and the development of efficient simplification strategies that can accelerate inference while preserving image quality.
Model Compression and Pruning. Mainstream compression techniques include knowledge distillation hinton2015distillation , quantization jacob2018quantization , and neural architecture search zoph2018nas . Unstructured pruning removes individual weights based on importance scores but typically results in irregular sparse patterns, which are difficult to accelerate on standard hardware without customized sparse kernels. In contrast, structured pruning li2017filter ; he2017channel simplifies the model by removing entire filters, channels, or blocks, offering better compatibility with existing hardware. However, few existing methods explicitly address the depth-wise redundancy of transformer-based generative models, especially in diffusion architectures, where deeply stacked structures are prevalent but costly.
Efficient Transformer and Block Simplification. Efficient Transformer variants choromanski2021performer ; tay2020long aim to reduce attention complexity through sparse or linear approximations, but primarily focus on token-level interactions rather than architectural depth. Other efforts simplify vision transformers by employing early exiting teerapittayanon2016branchynet ; elbayad2019depth , layer dropping fan2019layerdrop , or block merging kim2021learned , highlighting the unequal contribution of layers across depth. Recent works on structural adaptation hou2022dynamicvit ; liu2022depthprompting further demonstrate the importance of depth-aware design, revealing that static layer stacks are often suboptimal. Building on these insights, our method evaluates block-level importance via gradient-based and output-sensitivity analysis, enabling selective pruning or replacement of redundant layers in deep diffusion models while preserving generation quality.
@@
Figure 1: Overview of the FastFLUX Framework. (1) Importance Estimation: We compute the importance of each block and prioritize pruning of less important ones. (2) Block-wise Replacement with Linear Layers (BRLL): For each selected block, we replace its residual branch with a lightweight linear layer. The parameters of the linear layer are computed using the least squares method based on input-output pairs collected during forward inference. (3) Sandwich Training (ST): For non-overlapping cases, we train the replaced block along with its immediate neighbors enhanced by LoRA. For overlapping cases, we search upward or downward to locate the nearest unpruned blocks, apply LoRA to those, and jointly train all intermediate replaced blocks. Input-output data for each training step are re-collected by loading all previously trained sandwich structures. (4) LoRA Strategy: LoRA modules are selectively inserted into either single or double blocks during sandwich training to provide local supervision while keeping the rest of the model frozen. This framework enables efficient and structured pruning with minimal performance degradation.
@@ 3 FastFLUX: Block-wise Efficient Pruning with Sandwich Training
FastFLUX is designed to prune large-scale diffusion transformers in a lightweight and effective manner, improving inference efficiency while maintaining generation quality. In this section, we provide a complete description of our proposed pruning framework. In Section 3.1, we introduce Block-wise Replacement with Linear Layers (BRLL), which replace the residual branches of ResBlocks with simplified linear modules while preserving shortcut connections to retain key representational capacity. In Section 3.2, we present Sandwich Training (ST), a localized fine-tuning strategy that leverages LoRA-based supervision to compensate for potential performance loss after pruning, while maintaining high training efficiency. The overall workflow and implementation details of FastFLUX are illustrated in Figure 1 and Algorithm 1.
@@ 3.1 Block-wise Replacement with Linear Layer
To compress ultra-large models like FLUX and improve their inference efficiency, we propose an architecture-level pruning strategy called BRLL, which targets the internal residual structure of the model. We observe that FLUX consists of a large number of repeated ResBlocks with high parameter counts. Each ResBlock typically contains two branches: a computationally heavy residual branch and a lightweight shortcut branch. Our analysis indicates that the residual branch is the primary contributor to the model’s high computational cost and slow inference speed. Motivated by this, we propose to replace the residual branches with lightweight linear layers, while preserving the original shortcut connections during the replacement process. Although it is possible to replace the entire ResBlock with a single linear layer, we find that doing so leads to noticeable performance degradation, as the shortcut connection carries essential identity information. Therefore, our method selectively replaces only the residual branch, allowing the shortcut to preserve critical information flow. This distinguishes our approach from prior pruning methods that often disregard the functional asymmetry within ResBlocks. The method is illustrated in Figure 2.
Specifically, let X denote the input to a ResBlock, $f(X)$ the output of its residual branch, and $Y=f(X)+X$ the original output of the ResBlock. To reduce computation while preserving the residual structure, we replace the residual branch with a lightweight linear layer $g(X)=WX+b$, resulting in the modified output $\hat{Y}=g(X)+X$. To preserve the representational behavior of the original block, we minimize the approximation error between the original residual output and its linear approximation. This leads to the following objective: $\min_{W,b}||f(X)-g(X)||^{2}$. The optimal parameters $W^{*}$ of the linear layer can be obtained form using the least squares method, based on the collected input-output pairs of each ResBlock during forward inference:
$W^{*}=(f(X)-g(X))X^{T}(XX^{T})^{-1},W^{*}=[W;b].$ (1)
This procedure yields an initial estimation of the linear layer’s parameters, which can be further optimized during the subsequent training phase of our ST strategy described in Section 3.2.
Require : Pretrained FLUX model, replacement ratio $r\in[0,1]$, the number of blocks in the model $N$, the weight of FID in computing the importance score $\alpha$, the weight of CLIP score in computing the importance score $\beta$.
@@
                           $W^{*}=(f(X)-g(X))X^{T}(XX^{T})^{-1},W^{*}=[W;b]$
@@
      Replace the residual branch of $B_{k}$ with the linear layer using the weights $W*$;
@@ 3.2 Sandwich Training
To further enhance the generation quality of pruned models and bring them closer to the original model’s performance, we propose the Sandwich Training (ST) strategy. We observe that different blocks contribute unequally to the model’s performance, and prioritizing the pruning of less important blocks leads to more stable and effective compression. Therefore, we begin this section by introducing a method for Calculating Block Importance. Based on the computed importance scores, block pruning may lead to structural overlap between selected regions. To address this, our ST strategy includes dedicated mechanisms for Training Sandwich Blocks without Overlap and with Overlap, respectively. Finally, in Inference Reuse from High-Pruned Models, we present an empirical finding that transferring trained parameters from high-pruned models to lower-pruned configurations can further improve performance without additional training.
Calculating Block Importance. We observe that different blocks contribute unequally to the model’s generative performance (see Section 4.4 for more results). Based on this observation, we adopt an importance-aware pruning strategy that prioritizes the removal of less critical blocks. To quantitatively assess the importance of each block, we combine two commonly used metrics in text-to-image evaluation: FID heusel2018ganstrainedtimescaleupdate and CLIP Score radford2021learningtransferablevisualmodels . Specifically, for each ResBlock, we temporarily remove it from the model and generate a set of images. We then compute the FID and CLIP scores for each generated set and normalize both scores independently across all blocks. Since a lower FID indicates better quality, we transform it by taking $1-FID_{norm}$ so that higher values are consistently better across both metrics. The importance score $s_{i}$ for the i-th block is defined as:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the contribution of each metric. In our experiments, we set $\alpha\ =\beta\ =0.5$. A lower score $s_{i}$ implies that removing the block has a smaller impact on the overall generation quality, indicating that the block is less important and should be pruned earlier in the compression process. We sort all the blocks according to their importance scores in ascending order to conduct block replacement.
@@
Model Animation Concept-art Painting Photo Averaged Latency (ms)$\downarrow$
DALL·E miniramesh2021zeroshottexttoimagegeneration 26.10 25.56 25.56 26.12 25.83 -
@@
Model Single Two Counting Colors Position Attribute Overall Latency (ms)$\downarrow$
@@
DALL·E miniramesh2021zeroshottexttoimagegeneration 0.73 0.11 0.12 0.37 0.02 0.01 0.23 -
@@ 4.1 Implementation Details
Our base model is FLUX.1-dev, and all experiments are conducted using the fluxdev-controlnet-16k dataset, which consists of approximately 16.1k images. We sample 512 images to compute the parameters of the linear layers and another 512 images for training. To provide a comprehensive and holistic assessment of our method, we utilize a range of text-to-image evaluation benchmarks, including HPS v2 wu2023humanpreferencescorev2 and GENEVAL ghosh2023genevalobjectfocusedframeworkevaluating . All experiments are conducted in PyTorch and run on NVIDIA RTX 3090 GPUs.
@@ 4.2 Performance Comparison
Quantitative Results on Multiple Benchmarks. We evaluate the T2I generation performance of our model on two widely used benchmarks: HPS v2 (for assessing alignment with complex human preferences) and GENEVAL (for evaluating short-prompt generation). The results are shown in Table1 and Table2, respectively. We compare FastFLUX with the original FLUX.1-dev, L1-norm prunedliu2019rethinkingvaluenetworkpruning variants of FLUX, and several state-of-the-art T2I modelsrombach2022highresolutionimagesynthesislatent ; esser2024scaling ; ramesh2021zeroshottexttoimagegeneration . Since both FastFLUX and the L1-norm method are model compression techniques, we group models by similar inference latency for fair comparison. From the results, we observe that FastFLUX, when pruned by 10 percent, not only preserves generation quality but also outperforms the original FLUX.1-dev on the HPS v2 benchmark. Specifically, it achieves an improvement of 0.26 in score and ranks first among all evaluated T2I models. In addition, FastFLUX reduces inference latency by 7.48ms compared to the unpruned baseline. Within the same latency group, the L1-norm pruned model removes 12 percent of the parameters but performs 0.38 lower than FastFLUX on the HPS v2 metric. Moreover, L1-based pruning requires global fine-tuning to adapt to the irregular parameter structure, resulting in higher training costs. In contrast, FastFLUX requires only minimal localized adaptation. These observations are consistent across other latency groups and evaluation settings. Overall, FastFLUX significantly improves inference efficiency while maintaining high-quality image generation.
@@
Figure 7: HPS v2 scores under different numbers of layers involved in Sandwich Training. The x-axis indicates the number of layers with LoRA modules applied. The score increases significantly from 1 to 3 layers, then plateaus. A 3-layer structure is adopted as a trade-off between performance and efficiency.

### v2

@@ 1 Introduction
Recent advances in diffusion-based generative models zhang2024attentioncalibrationdisentangledtexttoimage; esser2024scaling; podell2024sdxl; sauer2023adversarialdiffusiondistillation; chen2023pixartalphafasttrainingdiffusion; zhang2023addingconditionalcontroltexttoimage have substantially improved the fidelity and controllability of text-to-image synthesis. Transformer-based architectures such as DiT peebles2022dit and FLUX flux2024 further enhance generation quality by stacking deep hierarchical blocks that increase model expressiveness and scalability. However, these designs also introduce significant computational overhead, including large memory consumption and high inference latency, which limits their suitability for real-time generation and deployment on resource-constrained hardware.
Although prior studies have investigated model compression through pruning castells2024ld; fang2023structural, knowledge distillation chen2025sana; meng2023distillation; salimans2022progressive, and quantization li2024svdquant; so2023temporal; he2023efficientdm, these approaches primarily focus on parameter-level optimization, largely overlooking the hierarchical structural organization of deep generative transformers. Specifically, hierarchical redundancy in diffusion-based generative models remains largely underexplored, despite being a significant source of inefficiency in modern deep transformer architectures. Furthermore, many existing techniques Shirkavand2025; fang2023structural require full-model fine-tuning after pruning or incur substantial performance degradation, which limits their practical applicability.
@@
In this work, we revisit the problem of hierarchical redundancy in large diffusion transformers and propose FastFLUX, a general and efficient framework for architecture-level compression. FastFLUX operates by selectively replacing the computationally heavy residual branches of FLUX blocks with lightweight linear alternatives, while preserving generation quality through a localized fine-tuning strategy. Specifically, we introduce Block-wise Replacement with Linear Layers (BRLL), which targets the high-cost residual pathways in ResBlocks. In FLUX, each ResBlock contains a heavy residual branch and a lightweight shortcut connection; the residual branch dominates both FLOPs and latency during inference. BRLL substitutes this branch with a simple linear layer while preserving the shortcut path, ensuring stable information flow. These linear layers can be initialized via least-squares estimation, providing a training-free initialization with controllable accuracy–efficiency trade-offs. To mitigate performance degradation after replacement, we propose a Sandwich Training (ST) strategy. Instead of full-model fine-tuning, ST performs localized adaptation by applying LoRA hu2021LoRAlowrankadaptationlarge modules to the neighboring unpruned layers of each replaced block, forming a small “sandwich” around the linear module. This local supervision enables the replaced branch to adapt effectively within its context, achieving stable and high-quality generation with minimal training overhead. Using these approaches, we prune 20% of the FLUX model while retaining high visual fidelity and significantly improving inference efficiency.
To sum up, our contributions are as follows:
@@
  • We introduce Block-wise Replacement with Linear Layers (BRLL), which replaces complex residual branches with linear layers. These replacements support training-free initialization via least-squares regression, with controllable performance degradation.
  • We propose Sandwich Training (ST), a localized fine-tuning strategy that applies LoRA to the adjacent unpruned layers of each replaced block to compensate for performance loss after pruning, enabling targeted adaptation with minimal training overhead.
@@ 2 Related Work
Diffusion-based Generative Models. Diffusion models have achieved remarkable progress across diverse domains cao2024survey, establishing foundations through discrete frameworks like DDPM ho2020ddpm and latent space formulations such as Latent Diffusion rombach2022ldm. Transformer-based architectures like DiT peebles2022dit demonstrate strong scalability, with recent models such as FLUX flux2024 further increasing depth to enhance generation quality. However, deeper hierarchical stacks introduce substantial computational and memory overhead, motivating the need for efficient compression strategies.
Model Quantization and Pruning. To reduce computational costs, diffusion models employ quantization and pruning techniques shen2025efficient. Post-training quantization methods like PTQ4DM shang2023post, Q-Diffusion li2023q, and PTQD he2023ptqd address timestep-specific calibration challenges, while quantization-aware training approaches such as TDQ so2023temporal and QALoRA he2023efficientdm learn optimized representations. Complementary pruning methods remove redundant weights via Taylor expansion fang2023structural, latent space guidance castells2024ld, or dynamic programming kim2024layermerge, achieving substantial size reduction with minimal quality loss.
Transformer Compression. Transformer compression tackles computational challenges through multiple strategies tang2024survey. Knowledge distillation transfers capabilities from large teachers to compact students sanh2019distilbert; touvron2021training, while quantization reduces precision via methods like FQ-ViT lin2021fq and PTQ-ViT liu2021post. Pruning techniques eliminate redundancy through head removal yu2023x or context selection anagnostidis2023dynamic. Architectural innovations target attention mechanisms ainslie2023gqa; hatamizadeh2023fastervit, feed-forward networks du2022glam; touvron2023llama, or propose alternative designs like Mamba gu2024mamba; zhu2024vision. While these methods address parameter and structural redundancy, explicit compression of depth-wise redundancy in diffusion transformers remains underexplored.
@@
Figure 1: Overview of the FastFLUX Framework. (1) Importance Estimation: Each block is ranked by estimated importance to determine the pruning order. (2) Block-wise Replacement with Linear Layers (BRLL): Selected blocks have their residual branches replaced with lightweight linear layers, whose parameters are initialized using a closed-form least-squares fit to the original block outputs. (3) Sandwich Training (ST): For non-overlapping cases, we train the replaced block along with its immediate neighbors enhanced by LoRA. For overlapping cases, we search upward or downward to locate the nearest unpruned blocks, apply LoRA to those, and jointly train all intermediate replaced blocks. Input-output data for each training step are re-collected by loading all previously trained sandwich structures. (4) LoRA Strategy: LoRA modules are selectively inserted into either single or double blocks during ST to provide local supervision while keeping the rest of the model frozen. This framework enables efficient and structured pruning with minimal performance degradation.
@@ 3 FastFLUX: Block-wise Efficient Pruning with Sandwich Training
FastFLUX is a lightweight and effective framework for pruning large-scale diffusion transformers, improving inference efficiency while preserving high-quality generation. We begin by introducing Block-wise Replacement with Linear Layers (BRLL), which substitutes the computationally heavy residual branches in ResBlocks with linear approximations while keeping the original shortcut pathways intact to maintain stable information flow. We then describe Sandwich Training (ST), a localized fine-tuning strategy that applies LoRA-based supervision to the neighboring unpruned layers of each replaced block, providing targeted adaptation with minimal training overhead. An overview of the full pipeline, together with implementation details, is provided in Figure 1 and Algorithm 1.
@@ 3.1 Block-wise Replacement with Linear Layer
To effectively compress ultra-large diffusion transformers such as FLUX, we propose a structural pruning strategy termed BRLL, which replaces the computationally expensive residual pathways of each block with lightweight linear modules. We observe that FLUX employs deep stacks of transformer-style ResBlocks, each of which consists of an expressive residual branch and a shortcut connection. Our profiling reveals that the residual branches dominate both parameter count and FLOPs, making them the primary bottleneck for inference. Motivated by this observation, BRLL substitutes only the residual branch with a lightweight linear approximation while preserving the original shortcut pathway. Unlike replacing the entire block, which results in severe degradation due to the loss of the identity pathway, selective replacement preserves stable information flow and significantly reduces computational cost. This distinguishes our approach from prior pruning methods that often disregard the functional asymmetry within ResBlocks. An overview of BRLL is provided in Figure 2.
[…]
$\tilde{W}^{*}=f(X)\tilde{X}^{T}(\tilde{X}\tilde{X}^{T})^{-1}.$ (1)
This procedure yields an efficient closed-form initialization of the linear layer’s parameters, which can be further optimized during the subsequent Sandwich Training phase.
Require : Pretrained FLUX model, replacement ratio $r\in[0,1]$, the number of blocks in the model NN, the weight of FID in computing the importance score $\alpha$, the weight of CLIP score in computing the importance score $\beta$.
@@
                       $\tilde{W}^{*}=f(X)\tilde{X}^{T}(\tilde{X}\tilde{X}^{T})^{-1},\tilde{X}=[X;\mathbf{1}],\tilde{W}=[W,b]$
@@
  Replace the residual branch of $B_{k}$ with the linear layer using the weights WW*;
@@ 3.2 Sandwich Training
To mitigate performance degradation after block replacement, we introduce Sandwich Training (ST), a localized fine-tuning strategy that adapts each replaced block using contextual supervision from its nearest unpruned neighbors. We observe that different blocks contribute unequally to the model’s performance, and prioritizing the pruning of less important blocks leads to more stable and effective compression. Therefore, we begin this section by introducing a method for Estimating Block Importance. Based on the computed importance scores, block pruning may lead to structural overlap between selected regions. To address this, our ST strategy includes dedicated mechanisms for Training Sandwich Blocks without Overlap and with Overlap, respectively. Finally, in Parameter Reuse across Pruning Ratios, we present an empirical finding that transferring trained parameters from high-pruned models to lower-pruned configurations can further improve performance without additional training.
Block Importance Estimation. We observe that different blocks contribute unequally to the model’s generative performance (see supplementary materials for more results). Based on this observation, we adopt an importance-aware pruning strategy that prioritizes the removal of less critical blocks. To quantitatively assess the importance of each block, we combine two commonly used metrics in text-to-image evaluation: FID heusel2018ganstrainedtimescaleupdate and CLIP Score radford2021learningtransferablevisualmodels. Specifically, for each ResBlock, we temporarily remove it from the model and generate a set of images. We then compute the FID and CLIP scores for each generated set and normalize both scores independently across all blocks. Since a lower FID indicates better quality, we transform it by taking $1-FID_{norm}$ so that higher values are consistently better across both metrics. The importance score $s_{i}$ for the i-th block is defined as:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the contribution of each metric. In our experiments, we set $\alpha\ =\beta\ =0.5$. A lower score $s_{i}$ implies that removing the block has a smaller impact on the overall generation quality, indicating that the block is less important and should be pruned earlier in the compression process. We therefore sort all blocks in ascending order of their importance scores and perform block replacement accordingly. Although Stable Flow StableFlow also employs layer-wise importance estimation, it relies solely on similarity-based metrics. In contrast, we incorporate both similarity and image quality assessments, enabling a more comprehensive evaluation of each block’s contribution to generation performance.
@@
[…]
DALL·E miniramesh2021zero 26.10 25.56 25.56 26.12 25.83 -
@@
[…]
@@
DALL·E miniramesh2021zero 0.73 0.11 0.12 0.37 0.02 0.01 0.23 -
@@ 4.1 Implementation Details
Our base model is FLUX.1-dev, and all experiments are conducted using the fluxdev-controlnet-16k dataset fluxdevcontrolnet16k, which consists of approximately 16.1k images. We sample 512 images to compute the parameters of the linear layers and another 512 images for training. Since we only need to fit the parameters of a single linear layer and a few LoRA modules at a time, only a small number of samples are required for training. We set the rank to 16 and the alpha to 8 of LoRA. To provide a comprehensive and holistic assessment of our method, we utilize a range of text-to-image evaluation benchmarks, including HPS v2 wu2023humanpreferencescorev2 and GENEVAL ghosh2023genevalobjectfocusedframeworkevaluating.All experiments are implemented in PyTorch and executed on two NVIDIA RTX 3090 GPUs using a block-swap mechanism for memory-efficient training.
@@ 4.2 Performance Comparison
Quantitative Results on Multiple Benchmarks. We evaluate the T2I generation performance of our model on two widely used benchmarks: HPS v2 (for assessing alignment with complex human preferences) and GENEVAL (for evaluating short-prompt generation). The results are shown in Table1 and Table2, respectively. We compare FastFLUX with the original FLUX.1-dev, L1-norm prunedwen2016learning; michel2019sixteen variants of FLUX, and several state-of-the-art T2I modelsramesh2021zero; rombach2022highresolutionimagesynthesislatent; rombach2022highresolutionimagesynthesislatent; esser2024scaling. Since both FastFLUX and the L1-norm method are model compression techniques, we group models by similar inference latency for fair comparison. From the results, we observe that FastFLUX, when pruned by 10 percent, not only preserves generation quality but also outperforms the original FLUX.1-dev on the HPS v2 benchmark. Specifically, it achieves an improvement of 0.26 in score and ranks first among all evaluated T2I models. In addition, FastFLUX reduces inference latency by 7.48ms compared to the unpruned baseline. Within the same latency group, the L1-norm pruned model removes 12 percent of the parameters but performs 0.38 lower than FastFLUX on the HPS v2 metric. Moreover, L1-based pruning requires global fine-tuning to adapt to the irregular parameter structure, resulting in higher training costs. In contrast, FastFLUX requires only minimal localized adaptation. These observations are consistent across other latency groups and evaluation settings. Overall, FastFLUX significantly improves inference efficiency while maintaining high-quality image generation.
@@
Figure 7: HPSv2 scores under different numbers of layers involved in ST. The x-axis indicates the total count of LoRA-enhanced layers. Performance improves substantially from 0 to 2 LoRA layers and saturates at 4, indicating a 2-layer structure provides a trade-off between performance and efficiency.